Week of 150713

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Monday

Attendance:

  • local: Andrea Sciabà, Christoph Wissing (CMS), Luca Canali (IT-DB), Maarten Litmaath (ALICE)
  • remote: Michael Ernst (BNL), Pavel Weber (KIT), Sang Un Ahn (KISTI), David Cameron (ATLAS), Christian (NDGF), Di Qing (TRIUMF), Lisa Giacchetti (FNAL), Tiju Idiculla (RAL)

Experiments round table:

  • ATLAS reports (raw view) -
    • LSF problem affecting T0 reported last week has gone
    • Still problems reading data from NIKHEF: GGUS:114431. Would be nice if a DPM/FTS expert could take a look.

  • CMS reports (raw view) -
    • Nothing really new since last meeting on Thursday
  • ALICE -
    • high activity

Sites / Services round table:

  • ASGC:
  • BNL: ntr
  • CNAF:
  • FNAL: ntr
  • GridPP:
  • IN2P3:
  • JINR:
  • KISTI: ntr
  • KIT:
  • NDGF: a full day of downtime this Thursday for an electrical power intervention in the Copenhagen site: all ALICE and ATLAS data there will be unavailable
  • NL-T1:
  • NRC-KI:
  • OSG:
  • PIC:
  • RAL: ntr
  • TRIUMF: ntr

  • CERN batch and grid services:
  • CERN storage services:
  • Databases: ntr
  • GGUS:
  • Grid Monitoring:
  • MW Officer:

AOB:

Thursday

Attendance:

  • local: A. Sciabà, D. Cameron (ATLAS), Nacho (Grid and batch services)
  • remote: Jeff Templon (NL-T1), Lisa Giacchetti (FNAL), Thomas Hartmann (KIT), Gareth Smith (RAL), Rob Quick (OSG), Christian (NDGF)

Experiments round table:

  • ALICE -
    • high activity
    • CERN: LSF cap removed on Tue July 14, thanks!
      • it used to be 15k and was very often reached during the last many weeks
      • the concurrent jobs are often 20k+ now
      • a top of 27k+ was reached for many hours!
    • CNAF: tape SE has been upgraded to Xrootd 4.1.3, thanks!
      • testing in progress
      • disk SE will follow

  • LHCb reports (raw view) -
    • Half of the jobs having problem to contact a vobox at CERN. Investigating.

Sites / Services round table:

  • ASGC:
  • BNL:
  • CNAF:
  • FNAL: ntr
  • GridPP:
  • IN2P3: ntr
  • JINR:
  • KISTI:
  • KIT: ntr
  • NDGF: downtime tomorrow for security update on dCache. It will be a rolling update and ATLAS and ALICE data will be temporarily unavailable.
  • NL-T1: ntr
  • NRC-KI:
  • OSG: ntr
  • PIC: ntr
  • RAL: ntr
  • TRIUMF: ntr

  • CERN batch and grid services: ntr
  • CERN storage services:
  • Databases:
  • GGUS:
  • Grid Monitoring:
  • MW Officer:

AOB:

Christian mentioned that NDGF is getting frequent tickets about missing files from the ATLAS shifters, which was due to the Rucio+FTS problem reported in the previous weeks and now solved. NDGF cannot do anything about that and it's not a site problem. David said that ATLAS is aware of that and will tell shifters not to open any ticket to sites for files from a list of known missing files.

Topic attachments
I Attachment History Action Size Date Who Comment
Unknown file formatpptx 2015-07-14.pptx r1 manage 2883.6 K 2015-07-13 - 15:50 MariaDimou GGUS Slides for the MB
Edit | Attach | Watch | Print version | History: r10 < r9 < r8 < r7 < r6 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r10 - 2015-07-17 - AndreaSciaba
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback